OpenStudy (anonymous):

What is the exact value of (sqrt) 99/55? Simplify if possible.

OpenStudy (anonymous):

start wtih \[\frac{99}{55}=\frac{9}{5}\]

OpenStudy (anonymous):

\[\sqrt{9}=3\] so one answer is \[\frac{3}{\sqrt{5}}\]

OpenStudy (anonymous):

3/sqrt5=3sqrt5/5

OpenStudy (anonymous):

I had an instructor that like to mark it wrong if i didnt rationalize :(